OpenStudy (darkigloo):

how do you simplify √((36-9x^2) / 4) ?

OpenStudy (darkigloo):

\[\sqrt{\frac{ 36-9x^2 }{ 4 }}\]

OpenStudy (darkigloo):

i used a calculator and it says \[\frac{ 3\sqrt{4 - x ^{2}} }{2 }\] but i don't know why.

OpenStudy (anonymous):

never use a calculator when it comes to this stuff they are dumb you just have to use your brain oh and @mathstudent55 I still need your help

OpenStudy (mathstudent55):

\(\LARGE\sqrt{\frac{ 36-9x^2 }{ 4 }}\) First, you can separate the numerator and the denominator by putting each one inside a root symbol. \(\LARGE \frac{ \sqrt{36-9x^2} }{\sqrt{ 4 }} \) The denominator is trivial: \(\LARGE \frac{ \sqrt{36-9x^2} }{2} \) Now work on the numerator by factoring a common factor. The GCF of the two terms is 9. \(\LARGE \frac{ \sqrt{9(4-x^2)} }{2} \) Now you can separate the two roots in the numerator: \(\LARGE \frac{ \sqrt{9}\sqrt{4-x^2} }{2} \) This now becomes: \(\LARGE \frac{ 3\sqrt{4-x^2} }{2} \) This is what you got.

OpenStudy (darkigloo):

:O ahh that makes sense. Thank you!